OpenStudy (anonymous):

A line contains the point (8, –5). If the slope of the line is 5/7, write the equation of the line using point-slope form.

pooja195 (pooja195):

Hi! Welcome to openstudy :) For this we would use the Point-Slope formula like this: \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Can you plug in the given values into the formula? :)